Goodman’s Bees Lodging Corporation had the following financial information for the past accounting period: - Current Liabilities = $286,000 - Long-term Liabilities = $167,500 - Total Fixed Assets = $310,000 - Current ratio = 1.45 Based on the financial information above, what is Goodman’s Bees Lodging Corporation’s debt to total assets ratio for the past accounting year? Group of answer choices .52 .46 .71 .63
Answer #1
Current ratio = Current assets / Current liabilities
1.45 = Current assets / 286000
Current assets = 1.45 * 286000 414700
Total debt = Current liabilities + Long term liabilities = 286000 + 167500 453500
Total assets = Total fixed assets + Current assets = 310000 + 414700 724700
Debt to total assets ratio = Total debt / Total assets = 453500 / 724700 0.63